# OneTrans: Unified Feature Interaction and Sequence Modeling with One Transformer in Industrial Recommender

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recommendation systems (RecSys) play a fundamental role in various information services, such as e-commerce (Zhou et al., 2018b; Feng et al., 2019), streaming media (Pancha et al., 2022; Chang et al., 2023; Xia et al., 2023) and social networks (Zhang et al., 2024). Industrial RecSys generally adopt a cascaded ranking architecture (Covington et al., 2016; Liu et al., 2017; Qin et al., 2022). First, a recall stage selects hundreds of candidates from billion-scale corpora (Zhu et al., 2018; Huang et al., 2024). Then, a ranking stage — often with coarse- and fine-ranking — scores each candidate and returns the top-$k$ items (Wang et al., 2021; Gui et al., 2023; Xia et al., 2023; Zhang et al., 2024; Zhu et al., 2025).
We focus on the ranking stage in this paper. For ranking, mainstream approaches iterate on two separate modules: (a) sequence modeling, which encodes user multi-behavior sequences into candidate-aware representations using local attention or Transformer encoders (Zhou et al., 2018b; Kang and McAuley, 2018; Sun et al., 2019; Chai et al., 2025), and (b) feature interaction, which learns high-order crosses among non-sequential features (e.g., user profile, item profile, and context) via factorization or explicit cross networks, or attention over feature groups (Guo et al., 2018; Wang et al., 2021; Gui et al., 2023; Zhu et al., 2025). As shown in Fig. 1(a), these approaches typically encode user behaviors into a compressed sequence representation, then concatenate it with non-sequential features and apply a feature-interaction module to learn higher-order interaction; we refer to this design as the encode-then-interaction pipeline.
@@ 4.1.1. Dataset
Table 2. Offline effectiveness (CTR/CVR) and efficiency; higher AUC/UAUC is better. Efficiency figures are collected from prior works, and dashes indicate unavailable. * indicates models deployed in our production lineage in chronological order: DCNv2+DIN $\rightarrow$ RankMixer+DIN $\rightarrow$ RankMixer+Transformer $\rightarrow$ OneTransS $\rightarrow$ OneTransL (default)
@@ 4.1.3. Baselines
We report two settings: OneTrans-S uses 6 stacked OneTrans blocks width d=256d{=}256, and H=4H{=}4 heads, targeting $\approx\!100$M parameters. OneTrans-L scales to 8 layers with width d=384d{=}384 (still H=4H{=}4). Inputs are processed through a unified tokenizer: multi-behavior sequences are fused in a timestamp-aware manner, while non-sequential features are tokenized via Auto-Split. The pyramid schedule linearly reduces tokens from 1190 to 12.
Optimization and infrastructure. We use a dual-optimizer strategy without weight decay: sparse embeddings are optimized with Adagrad ($\beta_{1}{=}0.1$, $\beta_{2}{=}1.0$), and dense parameters with RMSPropV2 (lr=0.005{=}0.005, momentum=0.99999{=}0.99999). The per-GPU batch size is set to 2048 during training, with gradient clipping thresholds of 90 for dense layers and 120 for sparse layers to ensure stable optimization. For online inference, we adopt a smaller batch size of 100 per GPU to balance throughput and latency. Training uses data-parallel all-reduce on 16 H100 GPUs.
@@ 4.3. RQ2: Design Choices via Ablation Study
We perform an ablation study of the proposed OneTransS model to quantify the contribution of key design choices. The complete results are summarized in Table 3. We evaluate the following variants: Input variants: i) Replacing the Auto-Split Tokenizer with a Group-wise Tokenizer (Row 1); ii) Using a timestamp-agnostic fusion strategy instead of the timestamp-aware sequence fusion (Row 2); iii) Removing [SEP] tokens in the timestamp-aware sequence fusion (Row 3); OneTrans block variants: i) Sharing a single set of Q/K/V and FFN parameters across all tokens, instead of assigning separate parameters to NS-tokens (Row 4); ii) Replacing causal attention with full attention (Row 5); iii) Disabling the pyramid stack by keeping the full token sequence at all layers (Row 6).
In summary, the ablations show that 1) Auto-Split Tokenizer provides a clear advantage over manually grouping non-sequential features into tokens, indicating that allowing the model to automatically build non-sequential tokens is more effective than relying on human-defined feature grouping; 2) Timestamp-aware fusion beats intent-based ordering when timestamps exist, suggesting that temporal ordering should be prioritized over event impact; 3) Under timestamp-agnostic fusion, learnable [SEP] tokens help the model separate sequences; 4) Assigning token-specific parameters to NS-tokens yields clear gains over sharing one set across all tokens, demonstrating that modeling non-sequential features with individualized projections enables better feature discrimination; 5) Causal and full attention achieve similar results, indicating that allowing tokens to attend to future positions is not crucial in this setting. Notably, we emphasize that full attention prohibits the use of standard optimizations such as KV caching; 6) Retaining the full token list at every layer provides no benefit: OneTrans effectively summarizes information into a small tail of tokens, so the pyramid design can safely prune queries to save computation.
@@ 4.4. RQ3: Systems Efficiency
Variant Training Inference
Runtime (ms) Memory (GB) Latency (p99; ms) Memory (GB)
Unoptimized OneTransS 407 53.13 54.00 1.70
+ Pyramid stack $28.7\%$ $42.6\%$ $8.4\%$ $6.9\%$
+ Cross-Request KV Caching $30.2\%$ $58.4\%$ $29.6\%$ $52.9\%$
+ FlashAttention $50.1\%$ $58.9\%$ $2.3\% 11$$1.6\%$
+ Mixed Precision with Recomputation $32.9\%$ $49.0\%$ $69.1\%$ $30.0\%$
To quantify the optimizations in Section 3.5, we ablate them on an unoptimized OneTransS baseline and report training/inference metrics in Table 5. The unoptimized OneTransS runs at 407 ms training runtime with 53.13 GB peak training memory, and 54.00 ms p99 inference latency with 1.70 GB inference memory, where p99 denotes the 99th-percentile (tail) latency—a standard SLO metric for high-availability online services. These differences reflect distinct operating conditions: offline training uses large per-device batches, while online inference distributes micro-batches across machines for stability. As shown in the table, 1) Pyramid stack yields substantial savings ($28.7\%$ training time, $42.6\%$ training memory, $8.4\%$ inference latency, $6.9\%$ inference memory) by compressing long behavioral histories into compact query sets; 2) Cross-request KV caching eliminates redundant sequence-side computation, reducing runtime/latency by ${\sim}30\%$ and memory by ${\sim}50\%$ in both training and serving; 3) FlashAttention primarily benefits training, reducing runtime by ${\sim}50\%$ and activation memory by ${\sim}58\%$. Inference gains are modest (${\sim}11$$12\%$ for latency and memory), as attention dominates training costs with larger batches and backpropagation; 4) Mixed precision with recomputation delivers the largest serving gains: p99 latency improves by ${\sim}69\%$ and inference memory by ${\sim}30\%$, as inference can operate end-to-end in low precision. By contrast, training must retain full-precision optimizer states and gradient accumulators; even so, training runtime and memory improve by $\sim$32% and $\sim$49%.
These results demonstrate the effectiveness of LLM optimizations for large-scale recommendation. Building on the ablations conducted on OneTransS, we scale up to OneTransL and show that with these techniques, OneTransL maintains online efficiency comparable to the much smaller DCNv2+DIN baseline (Table 4). This demonstrates again that reformulating RecSys into a unified Transformer backbone enables seamless adoption of LLM optimizations, unlocking effective scaling previously unattainable in traditional encode-then-interaction architectures.
4.5. RQ4: Scaling-Law Validation
[…]
Figure 3. Comparison of trade-off and scaling law.
@@ 4.6. RQ5: Online A/B Tests
We report user-level order/u and gmv/u as relative deltas ($\Delta\%$) versus the RankMixer+Transformer control with two-sided 95% CIs (user-level stratified bootstrap), and end-to-end latency—measured as the relative change in p99 per-impression time from request arrival to response emission ($\Delta\%$; lower is better). As shown in Table 6, OneTransL delivers consistent gains: in Feeds, $4.3510\%$ order/u, $5.6848\%$ gmv/u, and $3.91\%$ latency; in Mall, $2.5772\%$ order/u, $3.6696\%$ gmv/u, and $3.26\%$ latency — indicating the unified modeling framework lifts business metrics while reducing serving time relative to a strong non-unified baseline.
@@
Table 6. Online A/B results: OneTransL (treatment) vs. RankMixer+Transformer (control). Order/u and GMV/u are relative deltas (%). Latency is the relative end-to-end per-impression change $\Delta\%$ (lower is better). * denotes $p<0.05$, and ** for $p<0.01$
Scenario order/u gmv/u Latency ( $\Delta\%;p99$ ) $\downarrow$
Feeds $4.3510\%* +$$5.6848\%*$ $3.91\%$
Mall $2.5772\%** +$$3.6696\%*$ $3.26\%$

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systemsccs: Information systems Information retrievalccs: Information systems Recommender systems
@@ 1. Introduction
Recommendation systems (RecSys) play a fundamental role in various information services, such as e-commerce (Zhou et al., 2018b; Feng et al., 2019), streaming media (Pancha et al., 2022; Chang et al., 2023; Xia et al., 2023) and social networks (Zhang et al., 2024; Zhai et al., 2024). Industrial RecSys generally adopt a cascaded ranking architecture (Covington et al., 2016; Liu et al., 2017; Qin et al., 2022). First, a recall stage selects hundreds of candidates from billion-scale corpora (Zhu et al., 2018; Huang et al., 2024). Then, a ranking stage scores each candidate and returns the top-$k$ items (Wang et al., 2021; Gui et al., 2023; Xia et al., 2023; Zhang et al., 2024; Zhu et al., 2025). Deep Learning Recommendation Models (DLRMs) (Naumov et al., 2019) are widely adopted in the ranking stage of industrial recommenders.
We focus on the ranking stage in this paper, following the DLRM-style ranking paradigm. For ranking, mainstream approaches iterate on two separate modules: (a) sequence modeling, which encodes user multi-behavior sequences into candidate-aware representations using local attention or Transformer encoders (Zhou et al., 2018b; Kang and McAuley, 2018; Sun et al., 2019; Chai et al., 2025), and (b) feature interaction, which learns high-order crosses among non-sequential features (e.g., user profile, item profile, and context) via factorization, explicit cross networks, or attention over feature groups (Guo et al., 2018; Wang et al., 2021; Gui et al., 2023; Zhu et al., 2025). As shown in Fig. 1(a), these approaches typically encode user behaviors into a compressed sequence representation, then concatenate it with non-sequential features and apply a feature-interaction module to learn higher-order interaction; we refer to this design as the encode-then-interaction pipeline.
@@ 2. Related Work
Recent work on Generative Recommenders (GRs) frames recommendation as sequential transduction and proposes efficient long-context backbones such as HSTU (Zhai et al., 2024). This line is complementary to DLRMs that rely on rich non-sequential (NS) features.
@@ 4.1.1. Dataset
Table 2. Offline effectiveness (CTR/CVR) and efficiency; higher AUC/UAUC is better. * indicates models deployed in our production in chronological order: DCNv2+DIN $\rightarrow$ RankMixer+DIN $\rightarrow$ RankMixer+Transformer $\rightarrow$ OneTransS $\rightarrow$ OneTransL
@@ 4.1.3. Baselines
We report two settings: OneTransS uses 6 stacked OneTrans blocks width d=256d{=}256, and H=4H{=}4 heads, targeting $\approx\!100$M parameters. OneTransL scales to 8 layers with width d=384d{=}384.
Inputs are processed through a unified tokenizer (timestamp-aware fusion for multi-behavior sequences; Auto-Split for non-sequential features) and a heuristic pyramid schedule that, at each layer, linearly shrinks the number of sequential query tokens from 1190 to 12 (OneTransS) / from 1500 to 16 (OneTransL). Concretely, we linearly reduce the number of sequential query tokens across layers, rounding the token count at each layer to the nearest multiple of 32, and set the top layer to match the number of non-sequential tokens.
Optimization and infrastructure. We use a dual-optimizer strategy without weight decay: sparse embeddings are optimized with Adagrad ($\beta_{1}{=}0.1$, $\beta_{2}{=}1.0$), and dense parameters with RMSProp (lr=0.005{=}0.005, alpha=0.99999{=}0.99999,momentum=0{=}0). We apply stabilization techniques commonly used in large-scale Transformer training, including Pre-Norm (Xiong et al., 2020), and global grad-norm clipping (Shoeybi et al., 2019). The per-GPU batch size is set to 2048 during training, with gradient clipping thresholds of 90 for dense layers and 120 for sparse layers. For online inference, we adopt a smaller batch size of 100 per GPU to balance throughput and latency. Training uses data-parallel all-reduce on 16 H100 GPUs.
@@ 4.3. RQ2: Design Choices via Ablation Study
[…]
Figure 3. Comparison of trade-off and scaling law.
We perform an ablation study of the proposed OneTrans model to quantify the contribution of key design choices. The complete results are summarized in Table 3. We evaluate the following variants: Input variants: i) Replacing the Auto-Split Tokenizer with a Group-wise Tokenizer (Row 1); ii) Using a timestamp-agnostic fusion strategy instead of the timestamp-aware sequence fusion (Row 2); iii) Removing [SEP] tokens in the timestamp-aware sequence fusion (Row 3); OneTrans block variants: i) Sharing a single set of Q/K/V and FFN parameters across all tokens, instead of assigning separate parameters to NS-tokens (Row 4); ii) Replacing causal attention with full attention (Row 5); iii) Disabling the pyramid stack by keeping the full token sequence at all layers (Row 6).
In summary, the ablations show that 1) Auto-Split Tokenizer provides a clear advantage over manually grouping non-sequential features into tokens, indicating that allowing the model to automatically build non-sequential tokens is more effective than relying on human-defined feature grouping; 2) Timestamp-aware fusion beats intent-based ordering when timestamps exist, suggesting that temporal ordering should be prioritized over event impact; 3) Under timestamp-agnostic fusion, learnable [SEP] tokens help the model separate sequences; 4) Token-specific parameters for NS-tokens outperform a shared projection, enabling better feature discrimination; 5) Causal and full attention perform similarly, but full attention disables standard optimizations such as KV caching; 6) Keeping full-length tokens at all layers provides no benefit: OneTrans effectively summarizes information into a small tail, so the pyramid design can safely prune queries to save computation. Moreover, under a fixed TFLOPs budget, the pyramid design supports close to $1.75\times$ longer sequences than a full-length design, better exploiting gains from length extension.
@@ 4.4. RQ3: Systems Efficiency
Variant Training Inference
Runtime (ms) Memory (GB) Latency (p99; ms) Memory (GB)
Unoptimized OneTransS 407 53.13 54.00 1.70
+ Pyramid stack $28.7\%$ $42.6\%$ $8.4\%$ $6.9\%$
+ Cross-Request KV Caching $30.2\%$ $58.4\%$ $29.6\%$ $52.9\%$
+ FlashAttention $50.1\%$ $58.9\%$ $2.3\% 11$$1.6\%$
+ Mixed Precision with Recomputation $32.9\%$ $49.0\%$ $69.1\%$ $30.0\%$
To quantify the optimizations in Section 3.5, we ablate them on an unoptimized OneTransS baseline and report training/inference metrics in Table 4.
As shown in the table, (i) Pyramid stack reduces both training cost (runtime/memory) and serving overhead (p99 latency/memory) by pruning sequential query tokens; (ii) cross-request KV caching removes redundant sequence computation, consistently improving runtime/latency and memory in both training and serving; (iii) FlashAttention yields substantial training gains with modest serving improvements; and (iv) mixed precision with recomputation provides the largest serving gains (p99 latency and inference memory), while also improving training efficiency.
These results demonstrate the effectiveness of LLM optimizations for large-scale recommendation. Building on these results, we scale to OneTransL and show it maintains online efficiency comparable to the much smaller DCNv2+DIN baseline (Table 5), highlighting that a unified Transformer backbone enables direct adoption of LLM optimizations.
4.5. RQ4: Scaling-Law Validation
@@
While we can deploy OneTransL under strict online p99 latency constraints, scaling substantially beyond this regime remains constrained by online efficiency, and we leave further system–model co-optimizations to future work.
@@ 4.6. RQ5: Online A/B Tests
We report user-level click/u, order/u, and gmv/u as relative deltas ($\Delta\%$) versus the RankMixer+Transformer control with two-sided 95% CIs (user-level stratified bootstrap), and end-to-end latency, measured as the relative change in p99 per-impression time from request arrival to response emission ($\Delta\%$; lower is better). As shown in Table 6, OneTransL delivers consistent gains. In Feeds, it achieves $7.737\%$ click/u, $4.3510\%$ order/u, $5.6848\%$ gmv/u, and $3.91\%$ latency. In Mall, it achieves $5.143\%$ click/u, $2.5772\%$ order/u, $3.6696\%$ gmv/u, and $3.26\%$ latency. These results indicate that the unified modeling framework improves business metrics while reducing serving time relative to a strong non-unified baseline.
@@
Table 6. Online A/B results: OneTransL (treatment) vs. RankMixer+Transformer (control). Click/u, Order/u, GMV/u, are relative deltas (%). Latency is the relative end-to-end per-impression change $\Delta\%$ (lower is better). * denotes $p<0.05$, and ** for $p<0.01$
Scenario click/u order/u gmv/u Latency ( $p99$ ) $\downarrow$
Feeds $7.737\%** +$$4.351\%* +$$5.685\%*$ $3.91\%$
Mall $5.143\%** +$$2.577\%** +$$3.670\%*$ $3.26\%$